OpenStudy (anonymous):

what are the zeros of the function f(x)=(3x^2-3x-6)/(3x-6)

OpenStudy (anonymous):

–1 1 2 –2

OpenStudy (anonymous):

WILL GIVE MEDAL!!!!

OpenStudy (anonymous):

@thomaster @Juliette32801

OpenStudy (jdoe0001):

well, you have one there already, you just need to solve the quadratic factor there

OpenStudy (anonymous):

What do you mean?

OpenStudy (cwrw238):

3x^2 - 3x - 6 = 0 x^2 - x - 2 = 0 can u solve this?

OpenStudy (anonymous):

is it this: 2x^2-2x-4=0?

OpenStudy (cwrw238):

just solve x^2 - x - 2 = 0 by factoring then check if the roots are valid

OpenStudy (cwrw238):

you need to check because the equation is fractional - the denominator cannot be zero

OpenStudy (cwrw238):

(x -2)(x + 1) = 0 so x = 2 . -1 which one is valid?

OpenStudy (anonymous):

x=-1,2

OpenStudy (anonymous):

@cwrw238 so is it A OR C?

OpenStudy (anonymous):

@uri

OpenStudy (cwrw238):

it cannot be 2 because the denominator 3x - 6 gives 3(2) - 6 = 0